# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ B.2 Implicit bias in parameter space
$\min_{\tilde{\boldsymbol{a}}\in\mathbb{R}^{n}}\ D_{\Phi}(\tilde{\boldsymbol{a}},\boldsymbol{a}(0))\quad\mathrm{s.t.}\ \sum_{k\in[n]}\tilde{a}_{k}|w_{k}(0)x_{i}-b_{k}(0)|=y_{i},\ \forall i\in[m].$ (96)
@@
$g(x,\alpha)=\int\alpha(w,b)|wx-b|\ \mathrm{d}\mu.$
Note the above integral always annihilates the odd component of $\alpha$ since $|wx-b|$ is an even function of $(w,b)$ and $\mu$ is symmetric. Therefore, for networks with absolute value activation functions, we lose no generality to restrict the parameter space to the space of even and continuous functions:
@@
$\begin{dcases}g_{n}(x,\alpha)=\int\alpha(w,b)|wx-b|\ \mathrm{d}\mu_{n},\\ g(x,\alpha)=\int\alpha(w,b)|wx-b|\ \mathrm{d}\mu.\end{dcases}$ (98)
@@
Then given any $x\in\mathbb{R}$, with high probability over initialization, $\lim_{n\rightarrow\infty}|g_{n}(x,\bar{\alpha}_{n})-g(x,\bar{\alpha})|=0$.
@@
[Proof of Proposition 29] To ease the notation, for $k\in[n]$, let $z_{k}=(w_{k}(0),b_{k}(0))\in\mathbb{R}^{d+1}$ denote the input weight and bias associated with the $k$-th neuron at initialization. For $j\in[m]$, define $\sigma_{j}(z)=\sigma_{j}(w,b)=|wx_{j}-b|$. Let $\mathcal{Z}$ denote the random vector $(\mathcal{W},\mathcal{B})$, and $\tilde{y}_{i}=y_{i}-f(x_{i},\hat{\theta})$.
@@
The Lagrangian of problem 100 can be written as
$L(\alpha_{n},\mathbf{\lambda})=\int D_{\phi}(\alpha_{n}(w,b),0)\mathrm{d}\mu_{n}-\sum_{j=1}^{m}\lambda_{j}\Big{(}\int\alpha_{n}(z)\sigma_{j}(z)\mathrm{d}\mu_{n}-\tilde{y}_{i}\Big{)}.$
@@
Problem (100) is strictly convex in the quotient space of $\Theta_{\mathrm{Abs}}$ by the subspace of functions with zero $\ell_{2}(\mu_{n})$ semi-norm. Hence, $\bar{\alpha}_{n}$ is unique in the $\ell_{2}(\mu_{n})$ sense. Note that $(\phi^{\prime})^{-1}$ is injective, and that by Lemma 12, for large enough $n$ and with high probability, the matrix $J_{\boldsymbol{a}}\mathbf{f}(0)=(\sigma_{j}(z_{k}))_{j,k}$ is full row-rank. Therefore, the Lagrangian multipliers $\bar{\lambda}^{(n)}$ is uniquely determined by (101).
@@
Let $\mathbf{z}=(z_{1},\ldots,z_{n})^{T}$. We define a random $\mathbb{R}^{m}$-valued function as follows:
$\big{(}G_{n}(\lambda,\mathbf{z})\big{)}_{i}\triangleq\int(\phi^{\prime})^{-1}\Big{(}\phi^{\prime}(0)+\sum_{j=1}^{m}\lambda_{j}\sigma_{j}(z)\Big{)}\cdot\sigma_{i}(z)\ \mathrm{d}\mu_{n}-\tilde{y}_{i},\quad i=1,\ldots,m.$
@@
Since problem (99) is strictly convex in $\Theta_{\mathrm{Abs}}$, $\bar{\alpha}$ is unique. At the same time, note that $(\phi)^{-1}$ is injective, and that $\{\sigma_{j}\}_{j=1}^{m}$ are linearly independent in $L_{2}(\mu)$, as we have shown in the proof of Proposition 12. Hence, $\bar{\lambda}$ is the unique root of $G$.
@@
Therefore, $\langle\cdot,\cdot\rangle_{*}$ is a well-defined inner product on $L_{2}(\mu)$. Noting that $J_{\lambda}G(\bar{\lambda})$ can be the viewed as the Gram matrix of $\{\sigma_{l}\}_{l=1}^{m}$ with respect to $\langle\cdot,\cdot\rangle_{*}$, and that $\{\sigma_{l}\}_{l=1}^{m}$ are linearly independent in $L_{2}(\mu)$, we have $J_{\lambda}G(\bar{\lambda})$ is non-singular. Since $\bar{\lambda}^{(n)}$ and $\bar{\lambda}$ is the unique root of $G_{n}(\cdot)$ and $G(\cdot)$, respectively, by Lemma 30 we have
@@
$\begin{dcases}g_{n}(x,\bar{\alpha})=\int\bar{\alpha}(w,b)|wx-b|\ \mathrm{d}\mu_{n},\\ g(x,\bar{\alpha})=\int\bar{\alpha}(w,b)|wx-b|\ \mathrm{d}\mu.\end{dcases}$
@@
$\leq$ $\int\Big{|}\bar{\alpha}_{n}(w,b)-\bar{\alpha}(w,b)\Big{|}\sigma(wx-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\int\Big{|}(\phi^{\prime})^{-1}\Big{(}\phi^{\prime}(0)+\sum_{j=1}^{m}\bar{\lambda}^{(n)}_{j}\sigma_{j}(w,b)\Big{)}-(\phi^{\prime})^{-1}\Big{(}\phi^{\prime}(0)+\sum_{j=1}^{m}\bar{\lambda}_{j}\sigma_{j}(w,b)\Big{)}\Big{|}\sigma(wx-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\int L\Big{|}\sum_{j=1}^{m}\big{(}\bar{\lambda}^{(n)}_{j}-\bar{\lambda}_{j}\big{)}\sigma_{j}(w,b)\Big{|}\sigma(wx-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\int L\|\bar{\lambda}^{(n)}-\bar{\lambda}\|_{\infty}\sum_{j=1}^{m}\Big{(}\sigma(wx_{j}-b)\Big{)}\cdot\sigma(wx-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\Big{(}L\int\sigma(wx-b)\sum_{j=1}^{m}\sigma(wx_{j}-b)\ \mathrm{d}\mu_{n}\Big{)}\|\bar{\lambda}^{(n)}-\bar{\lambda}\|_{\infty}.$
@@
Hence $\int\sigma(wx-b)\sum_{j=1}^{m}\sigma(wx_{j}-b)\ \mathrm{d}\mu_{n}$ can be bounded by a finite number independent of $n$. Then by (104), we have
@@ B.3 Function description of the implicit bias
Now we study the solution to problem (99) in the function space
@@
It is clear that $s(x)\in\mathcal{F}_{\mathrm{Abs}}\subset\mathcal{F}_{\mathrm{Relu}}$. By Proposition 19, $s^{\prime\prime}$ vanishes when $x\neq\operatorname{supp}(p_{\mathcal{B}})$. For $x\in(\operatorname{supp}(p_{\mathcal{B}}))^{\mathrm{o}}$, similar to computation in (107) we also have $s^{\prime\prime}(x)=0$. Therefore, $s^{\prime}$ is a piece-wise affine function. According to Proposition 19, $s^{\prime}$ is continuous. Therefore, $s^{\prime}$ is a constant function. By what we just shown above, the only constant function in $\mathcal{F}_{\mathrm{Abs}}$ is the zero function. Therefore, $h(x)=g(x,\alpha_{h})$ for all $x\in\mathbb{R}$. Especially, we have that
@@
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\ \|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}$
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}$ that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
[Proof of Theorem 35] By Proposition 28, with sufficiently large $\beta$, Proposition 24 and 26 holds for $\phi=\rho_{\beta}$. Note in Proposition 29 we only require function $\phi$ to be $C^{3}$-smooth and to have positive second derivatives everywhere, which is satisfied by the hypentropy function. Also, notice Proposition 33 does not use information on the potential function. Therefore, Theorem 8 holds in this case and we conclude the proof. ∎
@@ Appendix C Discussion on parametrization and initialization
In initialization (2), we require the input weights to be sampled from the set $\{-1,1\}$. This choice aims to make the minimal representation cost problem (17) more tractable. Specifically, by this we have $\operatorname{supp}(\mu)=\{-1,1\}\times\mathbb{R}$ and the parameter space for infinitely wide ReLU networks, i.e. the space of uniformly continuous functions on $\operatorname{supp}(\mu)$, becomes much smaller than the case when input weights are sampled from a general sub-Gaussian distribution and $\operatorname{supp}(\mu)=\mathbb{R}\times\mathbb{R}$. We note this is a common setting in literature on representation cost for overparameterized networks (Savarese et al., 2019; Ongie et al., 2020). In Appendix E.2, we further discuss the challenge of solving the minimal representation cost problem for input weights sampled from general distributions.
@@
$f(x)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{k}\sigma(w_{k}x-b_{k})+d.$
Under NTK parametrization, a typical initialization scheme is to independently sample all parameters from a standard Gaussian distribution or, more generally, from a sub-Gaussian distribution. Note that, compared to standard parametrization in (1) and (2), the factor $1/\sqrt{n}$ appears before the sum, rather than as a scaling factor in the initialization distribution of the output weights. As a result, not only the output weights are normalized at initialization, their gradients are also normalized throughout training. It is known that, although gradient descent under both settings is in the kernel regime (lazy regime), but they correspond to training different kernel models and yield different implicit biases (Williams et al., 2019; Ghorbani et al., 2021). Specifically, noted by Jin & Montúfar (2023), under NTK parametrization one can not approximate the training dynamics of the full model by that of the model whose only output weights are trained. This implies the approach in this work, especially the formulation of problem 14, can not be directly applied to networks with NTK parametrization.
@@
$f(x,\theta)=\frac{1}{\sqrt{2}}\sum_{k=1}^{n}a_{k}[w_{k}x-b_{k}]_{+}-\frac{1}{\sqrt{2}}\sum_{k=1}^{n}a^{\prime}_{k}[w^{\prime}_{k}x-b^{\prime}_{k}]_{+}.$ (112)
@@
$f(x,(\boldsymbol{a},\boldsymbol{a}^{\prime}))=\frac{1}{\sqrt{2}}\sum_{k=1}^{n}(a_{k}-a^{\prime}_{k})\sigma\big{(}\hat{w}_{k}x-\hat{b}_{k}\big{)}.$ (113)
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))=-\eta(\nabla^{2}\Phi((\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))))^{-1}\cdot J_{(\boldsymbol{a},\boldsymbol{a}^{\prime})}\bar{\mathbf{f}}(t)\cdot(\bar{\mathbf{f}}(t)-\mathbf{y}),$ (114)
where $\bar{\mathbf{f}}(t)=(f(x_{1},(\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))),\ldots,f(x_{m},(\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))))^{T}\in\mathbb{R}^{m}$. Assume (114) is able to attain global optimality. Then by Theorem 10, the implicit bias is described by the following problem:
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(\boldsymbol{a}(t)-\boldsymbol{a}^{\prime}(t))=-\eta\Big{(}(\nabla^{2}\Phi(\boldsymbol{a}(t)))^{-1}+(\nabla^{2}\Phi(\boldsymbol{a}^{\prime}(t)))^{-1}\Big{)}J_{\boldsymbol{a}}\bar{\mathbf{f}}(0)(\bar{\mathbf{f}}(t)-\mathbf{y}).$ (116)
@@
If such a $\Psi$ exists, we have $\frac{\mathrm{d}}{\mathrm{dt}}(\nabla\Psi(\boldsymbol{a}(t)-\boldsymbol{a}^{\prime}(t)))=-\eta J_{\boldsymbol{a}}\bar{\mathbf{f}}(0)(\bar{\mathbf{f}}(t)-\mathbf{y}).$ It follows that $\boldsymbol{a}(\infty)-\boldsymbol{a}^{\prime}(\infty)$ is the solution to the following problem:

### v2

@@ B.2 Implicit bias in parameter space
$\min_{\tilde{\boldsymbol{a}}\in\mathbb{R}^{n}}\ D_{\Phi}(\tilde{\boldsymbol{a}},\boldsymbol{a}(0))\quad\mathrm{s.t.}\ \sum_{k\in[n]}\tilde{a}_{k}|w_{k}(0)^{T}x_{i}-b_{k}(0)|=y_{i},\ \forall i\in[m].$ (99)
@@
$g(x,\alpha)=\int\alpha(w,b)|w^{T}x-b|\ \mathrm{d}\mu.$
Note the above integral always annihilates the odd component of $\alpha$ since $|w^{T}x-b|$ is an even function of $(w,b)$ and $\mu$ is symmetric. Therefore, for networks with absolute value activation functions, we lose no generality to restrict the parameter space to the space of even and continuous functions:
@@
$\begin{dcases}g_{n}(x,\alpha)=\int\alpha(w,b)|w^{T}x-b|\ \mathrm{d}\mu_{n},\\ g(x,\alpha)=\int\alpha(w,b)|w^{T}x-b|\ \mathrm{d}\mu.\end{dcases}$ (101)
@@
Then given any $x\in\mathbb{R}^{d}$, with high probability over initialization, $\lim_{n\rightarrow\infty}|g_{n}(x,\bar{\alpha}_{n})-g(x,\bar{\alpha})|=0$.
@@
[Proof of Proposition 32] To ease the notation, for $k\in[n]$, let $z_{k}=(w_{k}(0),b_{k}(0))\in\mathbb{R}^{d+1}$ denote the input weight and bias associated with the $k$-th neuron at initialization. For $j\in[m]$, define $\sigma_{j}(z)=\sigma_{j}(w,b)=|w^{T}x_{j}-b|$. Let $\mathcal{Z}$ denote the random vector $(\mathcal{W},\mathcal{B})$, and $\tilde{y}_{i}=y_{i}-f(x_{i},\hat{\theta})$.
@@
The Lagrangian of problem 103 can be written as
$L(\alpha_{n},\boldsymbol{\lambda})=\int D_{\phi}(\alpha_{n}(w,b),0)\mathrm{d}\mu_{n}-\sum_{j=1}^{m}\lambda_{j}\Big{(}\int\alpha_{n}(z)\sigma_{j}(z)\mathrm{d}\mu_{n}-\tilde{y}_{i}\Big{)}.$
@@
Problem (103) is strictly convex in the quotient space of $\Theta_{\mathrm{Abs}}$ by the subspace of functions with zero $\ell_{2}(\mu_{n})$ semi-norm. Hence, $\bar{\alpha}_{n}$ is unique in the $\ell_{2}(\mu_{n})$ sense. Note that $(\phi^{\prime})^{-1}$ is injective, and that by Lemma 13, for large enough $n$ and with high probability, the matrix $J_{\boldsymbol{a}}\boldsymbol{f}(0)=(\sigma_{j}(z_{k}))_{j,k}$ is full row-rank. Therefore, the Lagrangian multipliers $\bar{\lambda}^{(n)}$ is uniquely determined by (104).
@@
Let $\boldsymbol{z}=(z_{1},\ldots,z_{n})^{T}$. We define a random $\mathbb{R}^{m}$-valued function as follows:
$\big{(}G_{n}(\lambda,\boldsymbol{z})\big{)}_{i}\triangleq\int(\phi^{\prime})^{-1}\Big{(}\phi^{\prime}(0)+\sum_{j=1}^{m}\lambda_{j}\sigma_{j}(z)\Big{)}\cdot\sigma_{i}(z)\ \mathrm{d}\mu_{n}-\tilde{y}_{i},\quad i=1,\ldots,m.$
@@
Since problem (102) is strictly convex in $\Theta_{\mathrm{Abs}}$, $\bar{\alpha}$ is unique. At the same time, note that $(\phi)^{-1}$ is injective, and that $\{\sigma_{j}\}_{j=1}^{m}$ are linearly independent in $L^{2}(\mu)$, as we have shown in the proof of Proposition 13. Hence, $\bar{\lambda}$ is the unique root of $G$.
@@
Therefore, $\langle\cdot,\cdot\rangle_{*}$ is a well-defined inner product on $L^{2}(\mu)$. Noting that $J_{\lambda}G(\bar{\lambda})$ can be the viewed as the Gram matrix of $\{\sigma_{l}\}_{l=1}^{m}$ with respect to $\langle\cdot,\cdot\rangle_{*}$, and that $\{\sigma_{l}\}_{l=1}^{m}$ are linearly independent in $L^{2}(\mu)$, we have $J_{\lambda}G(\bar{\lambda})$ is non-singular. Since $\bar{\lambda}^{(n)}$ and $\bar{\lambda}$ is the unique root of $G_{n}(\cdot)$ and $G(\cdot)$, respectively, by Lemma 33 we have
@@
$\begin{dcases}g_{n}(x,\bar{\alpha})=\int\bar{\alpha}(w,b)|w^{T}x-b|\ \mathrm{d}\mu_{n},\\ g(x,\bar{\alpha})=\int\bar{\alpha}(w,b)|w^{T}x-b|\ \mathrm{d}\mu.\end{dcases}$
@@
$\leq$ $\int\Big{|}\bar{\alpha}_{n}(w,b)-\bar{\alpha}(w,b)\Big{|}\sigma(w^{T}x-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\int\Big{|}(\phi^{\prime})^{-1}\Big{(}\phi^{\prime}(0)+\sum_{j=1}^{m}\bar{\lambda}^{(n)}_{j}\sigma_{j}(w,b)\Big{)}-(\phi^{\prime})^{-1}\Big{(}\phi^{\prime}(0)+\sum_{j=1}^{m}\bar{\lambda}_{j}\sigma_{j}(w,b)\Big{)}\Big{|}\sigma(w^{T}x-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\int L\Big{|}\sum_{j=1}^{m}\big{(}\bar{\lambda}^{(n)}_{j}-\bar{\lambda}_{j}\big{)}\sigma_{j}(w,b)\Big{|}\sigma(w^{T}x-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\int L\|\bar{\lambda}^{(n)}-\bar{\lambda}\|_{\infty}\sum_{j=1}^{m}\Big{(}\sigma(w^{T}x_{j}-b)\Big{)}\cdot\sigma(w^{T}x-b)\ \mathrm{d}\mu_{n}$
$\leq$ $\Big{(}L\int\sigma(w^{T}x-b)\sum_{j=1}^{m}\sigma(w^{T}x_{j}-b)\ \mathrm{d}\mu_{n}\Big{)}\|\bar{\lambda}^{(n)}-\bar{\lambda}\|_{\infty}.$
@@
Hence $\int\sigma(w^{T}x-b)\sum_{j=1}^{m}\sigma(w^{T}x_{j}-b)\ \mathrm{d}\mu_{n}$ can be bounded by a finite number independent of $n$. Then by (107), we have
@@ B.3 Function description of the implicit bias
Assume the input dimension is one, i.e., $d=1$. Now we study the solution to problem (102) in the function space
@@
It is clear that $s(x)\in\mathcal{F}_{\mathrm{Abs}}\subset\mathcal{F}_{\mathrm{Relu}}$. By Proposition 20, $s^{\prime\prime}$ vanishes when $x\neq\operatorname{supp}(p_{\mathcal{B}})$. For $x\in(\operatorname{supp}(p_{\mathcal{B}}))^{\mathrm{o}}$, similar to computation in (110) we also have $s^{\prime\prime}(x)=0$. Therefore, $s^{\prime}$ is a piece-wise affine function. According to Proposition 20, $s^{\prime}$ is continuous. Therefore, $s^{\prime}$ is a constant function. By what we just shown above, the only constant function in $\mathcal{F}_{\mathrm{Abs}}$ is the zero function. Therefore, $h(x)=g(x,\alpha_{h})$ for all $x\in\mathbb{R}$. In particular, we have that
@@
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\ \|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}$
Moreover, assuming univariate input data, i.e., $d=1$, and letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}$ that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
[Proof of Theorem 38] By Proposition 31, with sufficiently large $\beta$, Proposition 27 and 29 holds for $\phi=\rho_{\beta}$. Note in Proposition 32 we only require function $\phi$ to be $C^{3}$-smooth and to have positive second derivatives everywhere, which is satisfied by the hypentropy function. Also, notice Proposition 36 does not use information on the potential function. Therefore, Theorem 9 holds in this case and we conclude the proof. ∎
@@ Appendix C Discussion on parametrization and initialization
In initialization (2), we require the input weights to be sampled from the unit sphere $\mathbb{S}^{d-1}$. This choice aims to make the minimal representation cost problem (17) more tractable. Specifically, by this, for univariate input data we have $\operatorname{supp}(\mu)=\{-1,1\}\times\mathbb{R}$ and the parameter space for infinitely wide ReLU networks, i.e. the space of uniformly continuous functions on $\operatorname{supp}(\mu)$, becomes much smaller than the case when input weights are sampled from a general sub-Gaussian distribution and $\operatorname{supp}(\mu)=\mathbb{R}\times\mathbb{R}$. We note this is a common setting in literature on representation cost for overparameterized networks (Savarese et al., 2019; Ongie et al., 2020). In Appendix E.2, we further discuss the challenge of solving the minimal representation cost problem for input weights sampled from general distributions.
@@
$f(x)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{k}\sigma(w_{k}^{T}x-b_{k})+d.$
Under NTK parametrization, a typical initialization scheme is to independently sample all parameters from a standard Gaussian distribution or, more generally, from a sub-Gaussian distribution. Note that, compared to standard parametrization in (1) and (2), the factor $1/\sqrt{n}$ appears before the sum, rather than as a scaling factor in the initialization distribution of the output weights. As a result, not only the output weights are normalized at initialization, their gradients are also normalized throughout training. It is known that, although gradient descent under both settings is in the kernel regime (lazy regime), but they correspond to training different kernel models and yield different implicit biases (Williams et al., 2019; Ghorbani et al., 2021). Specifically, noted by Jin & Montúfar (2023), under NTK parametrization one can not approximate the training dynamics of the full model by that of the model whose only output weights are trained. This implies the approach in this work, in particular the formulation of problem 14, can not be directly applied to networks with NTK parametrization.
@@
$f(x,\theta)=\frac{1}{\sqrt{2}}\sum_{k=1}^{n}a_{k}[w_{k}^{T}x-b_{k}]_{+}-\frac{1}{\sqrt{2}}\sum_{k=1}^{n}a^{\prime}_{k}[(w^{\prime}_{k})^{T}x-b^{\prime}_{k}]_{+}.$ (115)
@@
$f(x,(\boldsymbol{a},\boldsymbol{a}^{\prime}))=\frac{1}{\sqrt{2}}\sum_{k=1}^{n}(a_{k}-a^{\prime}_{k})\sigma\big{(}\hat{w}_{k}^{T}x-\hat{b}_{k}\big{)}.$ (116)
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))=-\eta(\nabla^{2}\Phi((\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))))^{-1}\cdot J_{(\boldsymbol{a},\boldsymbol{a}^{\prime})}\bar{\boldsymbol{f}}(t)\cdot(\bar{\boldsymbol{f}}(t)-\boldsymbol{y}),$ (117)
where $\bar{\boldsymbol{f}}(t)=(f(x_{1},(\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))),\ldots,f(x_{m},(\boldsymbol{a}(t),\boldsymbol{a}^{\prime}(t))))^{T}\in\mathbb{R}^{m}$. Assume (117) is able to attain global optimality. Then by Theorem 11, the implicit bias is described by the following problem:
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(\boldsymbol{a}(t)-\boldsymbol{a}^{\prime}(t))=-\eta\Big{(}(\nabla^{2}\Phi(\boldsymbol{a}(t)))^{-1}+(\nabla^{2}\Phi(\boldsymbol{a}^{\prime}(t)))^{-1}\Big{)}J_{\boldsymbol{a}}\bar{\boldsymbol{f}}(0)(\bar{\boldsymbol{f}}(t)-\boldsymbol{y}).$ (119)
@@
If such a $\Psi$ exists, we have $\frac{\mathrm{d}}{\mathrm{dt}}(\nabla\Psi(\boldsymbol{a}(t)-\boldsymbol{a}^{\prime}(t)))=-\eta J_{\boldsymbol{a}}\bar{\boldsymbol{f}}(0)(\bar{\boldsymbol{f}}(t)-\boldsymbol{y}).$ It follows that $\boldsymbol{a}(\infty)-\boldsymbol{a}^{\prime}(\infty)$ is the solution to the following problem: